This problem does not come from a book, but I would appreciate it very much if you can solve it as I have a deadline tonight! Thanks so much~
Rewrite the equation below in general form Ax + By = C where A, B and C are the smallest integers possible and A > 0.
Enter the variable expression Ax + By separately from the constant term C.
y = 4/7(x) + 2/17

Answer by stanbon(48568)

y = (4/7)(x) + 2/17
-----------
Multiply thru by 7*17 = 119 to get:
119y = 4*17x + 2*7
119y = 68x + 14
68x+(-119)y = -14
---------
This is usually written:
68x - 119y = -14
================
Cheers,
Stan H.